A swimming pool is 24 feet wide by 45 feet long. George swims along the diagonal of the pool 5 times in one direction and 5 times in the other direction. What is the total number of feet that George swims?
Answer by rfer(16322)

c^2=45^2+24^2
c^2=2045+576
c^2=2601
c=51 ft
--------
10*51=510 ft total